1. Mr Grewal (m = 85 kg) and Ms Jang-Ing (m = 65 kg) are standing still on a 10 kg raft which is floating south down a river of 2.5 m/s. They both decide to jump off the raft at the same instant. Mr Grewal gives himself a velocity of 4.5 m/s [E] (with respect to the water) while Ms Jang-Ing has a velocity of 2.7 m/s [N 37° E] (with respect to the water). What is the velocity of the raft (with respect to the water) after they jump?


1
Expert's answer
2021-06-11T14:15:57-0400
"(85+65+10)(0,2)=(v_x,v_y)(10)\\+(85)(4.5,0)+(65)(2.7\cos{37},-2.7\sin{37})\\
(v_x,v_y)=(-52.3,42.6)\frac{m}{s}\\
v=\sqrt{52.3^2+42.6^2}=67.5\frac{m}{s}\\\theta=\arctan{\frac{42.6}{52.3}}=39\degree\text{ S of W}"